In two years, Edgard will be three-fourths as old as his sister, Edna. Two years ago, he was two-thirds as old as she. How old are they?: In two years, Edgard will be three-fourths as old as his sister, Edna. Two years ago, he was two-thirds as old as she. How old are they?
Answer by checkley77(3423)

x+2=3y/4 or x=3y/4-2
x-2=2y/3
(3y/4-2)-2=2y/3
(3y-2*4)/4-2=2y/3
(3y-8)/4-2=2y/3
(3y-8-2*4)/4=2y/3
(3y-8-8)/4=2y/3
(3y-16)/4=2y/3 cross multiply.
3(3y-16)=4*2y
9y-48=8y
9y-8y=48
y=48 age of Edgard now.
x=3*48/4-2
x=3*12-2
x=36-2
x=34 age of Edna now.
Proof:
34-2=2*48/3
32=96/3
32=32